the sum of two integers is 5. their difference is 13. find the two integers
Found 3 solutions by nerdybill, Maths68, xdragonfight:
Answer by nerdybill(7384)

the sum of two integers is 5. their difference is 13. find the two integers
.
Let x = one integer
and y = other integer
then
x+y = 5
x-y = 13
.
add both equations together:
x+y = 5
x-y = 13
------------
2x = 18
x = 9
.
substitute above into:
x+y = 5
9+y = 5
y = -4
.
answer 9 and -4

Answer by Maths68(1474)

Let x and y are two integers.
x+y=5.......(1)
x-y=13......(2)
Add (1) and (2)
x+y=5.......(1)
x-y=13......(2)
-------
2x=18
2x/2=18/2
x=9
Put the value of x in (1)
x+y=5
9+y=5
y=5-9
y=-4
Integers are -4 and 9


Check
========
their sum = 5
9+(-4)=5
9-4=5
5=5
Their difference = 13
9-(-4)=13
9+4=13
13=13

Answer by xdragonfight(116)

x + y = 5
x - y = 13